# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 4.1.2. Other EEG-based ADA systems
To extract depression-related features from EEG signals, Mohammadi et al. (Mohammadi et al., 2015) utilises Linear Discriminant Analysis (LDA) and Genetic Algorithm for the feature extraction. These features are then fed to a decision tree (DT) to predict the depression score. While hand-crafted features for EEG-based ADA are tedious, time consuming (Ay et al., 2019) and require domain knowledge, recent studies (Zhu et al., 2020; Dang et al., 2020; Wan et al., 2020; Tian et al., 2024; Ying et al., 2024; Kwok et al., 2025) also deep learn depression-specific features from EEG signals for ADA. For example, Seal et al. (Seal et al., 2021) propose a CNN to process the pictorial representation of the given EEG signal for ADA. Betul et al. (Ay et al., 2019) combine CNN and Long-short-term-memory (LSTM) network to process EEG signals obtained from left and right hemispheres of human brain for assessing depression level. Similar CNN-LSTM architecture also has been proposed by (Song et al., 2022c) for EEG-based ADA. Aiming at real-time EEG-based ADA, Sharma et al. (Sharma et al., 2023) first compute a spectrogram image from the target EEG signal using STFT, and then feed it to a CNN-LSTM or CNN-Gated Recurrent Units (GRU) network for classification. A couple of transformer networks have been extended for EEG-based ADA, including attention-based gated recurrent units time-series transformer model (Tigga and Garg, 2022), Decentralised-Centralised Structure Transformer (Shao et al., 2024b), and Graph Convolutional Transformer (Wang et al., 2024b). These approaches directly applied existing transformer architecture for EEG-based ADA but did not investigate how to design depression-specific modules.
4.2. MRI-based depression assessment
@@
The common challenges in MRI-based ADA includes how to model spatio-temporal nature of the MRI data, and identifying the conductivity patterns/regions relevant to depression within MRI data. To model spatio-temporal cues, CNN models (Mousavian et al., 2020; Chen et al., 2021b; Li et al., 2024a) have been introduced to learn depression-specific spatio-temporal cues from MRI images, e.g., Pominova et al. (Pominova et al., 2018) employ multiple deep voxelwise networks to assess depression on structural (3D) and functional (4D) MRI data, each of which is built on several residual volumetric convolutional modules. Alternatively, an unsupervised discrepancy-based cross-domain fMRI adaptation framework is proposed (Fang et al., 2023b), which applies attention-guided GCN to model spatio-temporal fMRI patterns for ADA. Similarly, Yao et al. (Yao et al., 2020) first segment each region-of-interest (ROI)-based fMRI image series and then propose a Temporal-Adaptive GCN to model depression-related spatio-temporal cues. For the latter challenge, Gao et al. (Gao et al., 2023) first define a heatmap of conductivity patterns indicating positive influences within the brain, which then guides a lightweight 3D CNN in learning depression cues from 3D Gray Matter data recorded by fMRI, where patterns reflected by prefrontal cortex, insula, superior temporal cortex, and cingulate cortex have been identified to be linked with depression. Qin et al. (Qin et al., 2022) employ GCN to process rsfMRI for depression assessment, showing that the most salient regions related to depression are within the default mode, fronto-parietal, and cingulo-opercular networks, while Nodal topologies of the left inferior parietal lobule is also associated with depression symptoms. We discuss other MRI-based ADA methods in the subsection below.
4.2.2. Other MRI-based ADA systems.
@@
Gao et al. (Gao et al., 2018) review 63 approaches and conclude that a common pipeline for MRI-based ADA approaches typically consists of three stages: (i) MRI data pre-processing, (ii) depression feature extraction, and (iii) ADA model training/inference. A comparative study (Winter et al., 2024) systematically evaluates the feasibility of applying traditional ML models to analyse biomarkers provided by sMRI, fMRI and DTI images for ADA, where various traditional ML models (e.g., SVM, KNN, RF, LR, naive Bayers and boosting classifier) were employed, with diagnostic classification accuracy ranged between 48.1% and 62.0%. Recently, spatial GCN (Dai et al., 2023; Jun et al., 2020; Zhu et al., 2023) and spatio-temporal GCN (Kong et al., 2021) have been found to be particularly effective for MRI-based ADA. Ktena et al. (Ktena et al., 2018) propose a Siamese GCN that adopts spectral graph convolutions to process fMRI data for ADA. It encodes each fMRI time series as a graph with topology decided by Pearson Correlation Coefficient (PCC) computed between time-series channels, where a novel graph similarity metric is introduced to measure the similarity between each pair of fMRI graphs during training. Meanwhile, Mousavian et al. (Mousavian et al., 2021) propose a unified framework for sMRI and rsfMRI-based ADA, including data preparation (e.g., removing backgrounds, normalisation and data resize), 3D region definition within the MRI image, whole-brain matter similarity computation based on pre-defined 3D regions, spectral domain feature extraction, Wilcoxon rank-sum test-based feature selection, and naive Bayers-based depression classification.
4.3. Face/head-based depression assessment
@@
Since human facial/head behaviours provide informative depression biomarkers (discussed in Sec. 2.2.4) while recent advanced DL models can accurately and automatically recognise various facial/head behaviours (e.g., AUs, head movements and facial expressions), automatic visual behaviour-based depression assessment is becoming a prevalent research topic in the past decade. Existing facial/head behaviour-based ADA approaches can be categorised into three types: (i) face image-based approaches which assess depression merely based on a static facial display (e.g., a video frame) (Valstar et al., 2013, 2014); (ii) video-based approaches which assess depression based on spatio-temporal facial behaviours (Xu et al., 2024; de Melo et al., 2021; Harati et al., 2019; Tao et al., 2019; He et al., 2022a; Du et al., 2019); and (iii) facial attribute-based approaches which assess depression based on pre-extracted anonymous facial descriptors (e.g., facial landmarks (Pan et al., 2023; Niu et al., 2024a), AUs (Song et al., 2018, 2022a; Jaiswal et al., 2019), head poses (Valstar et al., 2016; Ringeval et al., 2017), gazes/blinks (Yang et al., 2022; Alghowinem et al., 2015) or frame-level deep facial features (Ringeval et al., 2019)).
4.3.1. Challenges.
@@
(i) a key challenge for face-based ADA is that depression is more reliably to be reflected by spatio-temporal facial behaviours rather than a static image, i.e., depressed and non-depressed individuals can express same facial displays in a single frame. However, some ADA approaches still model frame-level depression-related facial cues despite the availability of face videos, which frequently aggregate frame-level predictions via average or majority voting for video-level assessment. For example, the baselines and participants of AVEC 2013 (Valstar et al., 2013) and AVEC 2014 challenges (Valstar et al., 2014), generally extract different hand-crafted image-level descriptors (e.g., Histogram of oriented gradients (HOG) (Dalal and Triggs, 2005) and Local Phase Quantisation (LPQ) (Ojansivu and Heikkilä, 2008)), and then feed them to SVM Regression (SVR) for assessing frame-level depression severity. Following a similar pipeline, several early studies also extract Gabor wavelet features (Girard et al., 2013; Jan et al., 2014) from static face images, with KNN as the depression predictor. However, such approaches re-use the video-level label as the frame-level labels to train models, and thus may cause ‘ill-posed problems’ (i.e., similar static facial expression inputs paired with different depression labels), making it impossible to learn good hypothesis. To address this issue, spatio-temporal modelling strategies have been particularly explored for ADA (please refer to Sec. 4.3.3 for details); (ii) while depression is associated with long-term and multi-scale facial behaviours (Song et al., 2022a), most existing spatio-temporal ADA methods only learn depression cues from short segments, ignoring crucial long-term (i.e., video-level) but variable-length spatio-temporal depression cues at the feature extraction level (Al Jazaery and Guo, 2018; Pan et al., 2024b). This is partially addressed by video-level modelling ADA approaches. Song et al. (Song et al., 2022a, 2018) propose a spectral encoding strategy which treats the facial primitives as a multi-channel time-series, applying Fourier Transform to generate a raw spectral representation, and then normalises its size through frequency alignment to standardise all frame-level features as a fixed-size video-level representation. Building on this, Chen et al. (Chen et al., 2022) extend the method by incorporating a Neural Architecture Search (NAS) in order to identify optimal networks for inferring depression status from the video-level facial primitive spectral representations. Another study (Chen et al., 2021a) investigates a sequential chained-fusion strategy to extract complementary depression patterns contained in both facial display and facial dynamics for the video-level depression assessment. Melo et al. (de Melo et al., 2021) propose a maximization branch to learn multi-scale smooth facial variations and a difference branch for encoding multi-scale facial dynamics from the video-level image sequence, both of which are fused for video-level ADA; (iii) the third limitation is that most existing ADA methods were devoted to assessing adults’ depression status. To bridge this gap, Abbasi et al. (Abbasi et al., 2022) construct a video-level spatio-temporal graph for assessing children’s depression status; (iv) another challenge for inferring depression from facial/head behaviours is that these behavuours are more prone to data privacy and ethical issues. Thus, some ADA studies (Song et al., 2022a; Niu et al., 2024a; Valstar et al., 2016, 2013) were developed to infer depression from anonymous facial attributes (discussed in Sec. 4.3.4). Alternatively, Pan et al. (Pan et al., 2024a) design a privacy preservation ADA system which records de-identified facial images, while retaining not only depression-relevant features but also emotion-related cues for face image-level ADA; and (v) face/head data can also suffer from noisy and ambiguous depression annotations. He et al. (He et al., 2022c) proposes three modules (self-attention, square ranking regularization and re-label modules) to reduce the impact of noisy and ambiguous depression annotations. To deal with missing faces, face detection errors, frame loss and incorrect labels, Song et al. (Song et al., 2024) introduces a generic loss function relaxation strategy to mitigate abnormal loss values for various loss functions during video-based ADA model training. We further discuss other existing face/head behaviour-based ADA approaches in the following sub-paragraphs.
4.3.2. Face image-based approaches.
@@
while depressed and non-depressed individuals can express same static facial displays (e.g., neutral face), some approaches still attempt to assess depression from a single face image (e.g., facial expression and pupil (Liu et al., 2024; Li et al., 2024b)). Melo et al. (De Melo et al., 2019) propose a distribution learning strategy to train a ResNet for face image-based ADA. Liu et al. (Liu et al., 2022) pre-train a CNN network on a facial expression dataset and then linearly superimpose the entropy of the detected expression into depression severity. Zhao et al. (Zhao et al., 2023) propose a frequency attention strategy to identify depression from static facial displays. Li et al. (Li et al., 2024b) integrate a dual-scale convolution and a channel-wise adaptive attention to a standard CNN network, aiming to infer depression from a facial expression. Zhang et al. (Zhang et al., 2023b) propose a two-stream deep network LKCT. A decoupled knowledge distillation module then distills the LKCT into the lightweight network for face image-based ADA. Similarly, a standard frame-level DL-based ADA pipeline also learns depression cues from individual face frames and aggregates predictions or features for video-level assessment. Zhou et al. (Zhou et al., 2018) propose a DepressNet consisting of two convolution layers and four bottleneck residual convolution blocks to identify depression-informative facial regions from each face frame for ADA. Shang et al. (Shang et al., 2021) propose to extract each facial image with local textures in the quaternion domain to preserve the dependence of color channels, while deep learning global facial features via a simple CNN. Then, it fuses hand-crafted local features with the deep-learned global cues for frame-level depression assessment. To model multi-scale spatial facial cues for ADA, He et al. (He et al., 2021) propose local and global attention blocks to capture depression-related facial features at multiple spatial scales, which are then combined via a Weighted Spatial Pyramid Pooling for image-level ADA.
4.3.3. Face video-based approaches.
@@
Facial temporal dynamics contain crucial biomarkers (discussed in Sec. 2.2.4) for inferring depression. Early face video-based ADA approaches frequently summarise video-level spatio-temporal facial behaviours by computing either the video-level statistics of hand-crafted frame/segment-level facial features (e.g., the major facial expressions, variance of facial expressions (Scherer et al., 2013b)) or differences/dynamics between facial frames such as velocity of angular (Girard et al., 2014), facial appearance eigenviectors (Cohn et al., 2009), three orthogonal planes-based dynamic representations (e.g., LBP-TOP, LGBP-TOP, LPQ-TOP) (Joshi et al., 2013b; Gupta et al., 2014; Kaya et al., 2014; He et al., 2018; Dhall and Goecke, 2015), Motion History Histogram (MHH) (Meng et al., 2013), and pixel differences between adjacent frames (Kächele et al., 2014). For instance, Pampouchidou et al. (Pampouchidou et al., 2015) extract three types of features, including appearance-based features (LBP and LBP-TOP), model-based features (active appearance models (AAMs)), and motion-based features (i.e., frame differencing, feature point tracking and optical flow), from every frame/short segment of the given face video, and then feed their statistics to a SVR for ADA. He et al. (He et al., 2018) extend the LBP-TOP feature to MRLBP-TOP for extracting short-term dynamics and then apply Fisher Vector to aggregate them for video-level ADA. SVR then integrates decisions across dimensions for video-level ADA. Instead of directly learning depression cues from face videos, Casado et al.(Casado et al., 2023) calculate over 60 video-level statistical, geometrical, and physiological features from the remote photoplethysmography signals obtained from each face video for ADA. Yang et al. (Yang et al., 2023a) proposes a Trial Selection Tensor Canonical Correlation Analysis to fuse spatio-temporal relevance between facial dynamic features and pupil diameter features for ADA due to their complementary affective nature.
@@
A prototypical DL solution for face video-based ADA is through learning frame-level facial features/predictions and aggregating them as a video-level depression representation. Uddin et al. (Uddin et al., 2020) apply LSTM to model the facial dynamics from CNN-learned frame-level depression features. Xu et al. (Xu et al., 2024) first encode a multi-scale facial spatial representation from every frame of the target video, and then integrate all frame-level multi-scale spatial features as a single spectral graph representation for video-level ADA in an effort to learn a richer and more condensed latent representation. Pan et al. (Pan et al., 2024b) integrate both spatial and temporal attention mechanisms to 3D CNNs, and then combine all segment-level depression predictions as the video-level prediction. Similarly, Al et al. (Al Jazaery and Guo, 2018) utilise a C3D network to extract short segment-level facial dynamics and then averages all segment-level predictions for video-level ADA. Li et al. (Li et al., 2024c) and Melo et al. (de Melo et al., 2019) propose two-stream networks that jointly learns facial image-level depression cues and video-level facial dynamics, both of which are concatenated via a Multi-Layer Perceptron (MLP) for the final video-level ADA. Zhu et al. (Zhu et al., 2017) propose two-stream networks that jointly learns facial appearance features and short-term facial dynamics via optical flow. Zhang et al. (Zhang et al., 2023a) propose a lightweight CNN incorporating temporal difference for short-term facial behaviours, followed by an attention for long-term facial behaviour modelling, which are fused via Interactive Multi-head Attention for video-level ADA. He et al. (He et al., 2024) employ a video Swin-Transformer to learn spatio-temporal and fine-grained depression-specific features, integrating them via a Depression Channel Attention Fusion (DCAF) module for video-level ADA. Xie et al. (Xie et al., 2021) design an end-to-end ADA framework for variable-length videos, which leverages a 3D CNN for local temporal pattern extraction and a Redundancy-Aware Self-Attention (RAS) scheme for global feature aggregation. Moreno et al. (Moreno et al., 2023) pre-train a CNN on action recognition datasets and fine-tune it on ADA face videos, enabling analysis of frame-level regions and short-term temporal expressions.
4.3.4. Facial behaviour primitive-based approaches.

### v2

@@ 4.1.2 Other EEG-based ADA systems
To extract depression-related features from EEG signals, Mohammadi et al. [148] utilises Linear Discriminant Analysis (LDA) and Genetic Algorithm for the feature extraction. These features are then fed to a decision tree (DT) to predict the depression score. While hand-crafted features for EEG-based ADA are tedious, time consuming [153] and require domain knowledge, recent studies [149, 151, 152, 156, 157, 277] also deep learn depression-specific features from EEG signals. For example, Seal et al. [51] propose a CNN to process the pictorial representation of the given EEG signal for ADA. Betul et al. [153] combine CNN and Long-short-term-memory (LSTM) network to process EEG signals obtained from left and right hemispheres of human brain for assessing depression level. Similar CNN-LSTM architecture also has been proposed by [154] for EEG-based ADA. Aiming at real-time EEG-based ADA, Sharma et al. [52] first compute a spectrogram image from the target EEG signal using STFT, and then feed it to a CNN-LSTM or CNN-Gated Recurrent Units (GRU) network for classification. A couple of transformer networks have been extended for EEG-based ADA, including attention-based gated recurrent units time-series transformer model [155], Decentralised-Centralised Structure Transformer [53], and Graph Convolutional Transformer [54]. These approaches directly applied existing transformer architecture for EEG-based ADA but did not investigate how to design depression-specific modules.
4.2 MRI-based depression assessment
@@ 4.2.1 Challenges and solutions
The common challenges in MRI-based ADA includes how to model spatio-temporal nature of the MRI data, and identifying the conductivity patterns/regions relevant to depression within MRI data. To model spatio-temporal cues, CNN models [169, 46, 168] have been introduced to learn depression-specific spatio-temporal cues from MRI images, e.g., Pominova et al. [161] employ multiple deep voxelwise networks to assess depression on structural (3D) and functional (4D) MRI data, each of which is built on several residual volumetric convolutional modules. Alternatively, an unsupervised discrepancy-based cross-domain fMRI adaptation framework is proposed [165], which applies attention-guided GCN to model spatio-temporal fMRI patterns for ADA. Similarly, Yao et al. [160] first segment each region-of-interest (ROI)-based fMRI image series and then propose a Temporal-Adaptive GCN to model depression-related spatio-temporal cues. For the latter challenge, Gao et al. [47] first define a heatmap of conductivity patterns indicating positive influences within the brain, which then guides a lightweight 3D CNN in learning depression cues from 3D Gray Matter data recorded by fMRI, where patterns reflected by prefrontal cortex, insula, superior temporal cortex, and cingulate cortex have been identified to be linked with depression. Qin et al. [166] employ GCN to process rsfMRI for depression assessment, showing that the most salient regions related to depression are within the default mode, fronto-parietal, and cingulo-opercular networks, while Nodal topologies of the left inferior parietal lobule is also associated with depression symptoms. We discuss other MRI-based ADA methods in the subsection below.
4.2.2 Other MRI-based ADA systems
@@
Gao et al. [71] review 63 approaches and conclude that a common pipeline for MRI-based ADA approaches typically consists of three stages: (i) MRI data pre-processing, (ii) depression feature extraction, and (iii) ADA model training/inference. A comparative study [159] systematically evaluates the feasibility of applying traditional ML models to analyse biomarkers provided by sMRI, fMRI and DTI images for ADA, where various traditional ML models (e.g., SVM, KNN, RF, LR, naive Bayers and boosting classifier) were employed, with diagnostic classification accuracy ranged between 48.1% and 62.0%. Recently, spatial GCN [57, 163, 58] and spatio-temporal GCN [170] have been found to be particularly effective for MRI-based ADA. Ktena et al. [164] propose a Siamese GCN that adopts spectral graph convolutions to process fMRI data for ADA. It encodes each fMRI time series as a graph with topology decided by Pearson Correlation Coefficient (PCC) computed between time-series channels, where a novel graph similarity metric is introduced to measure the similarity between each pair of fMRI graphs during training. Meanwhile, Mousavian et al. [48] propose a unified framework for sMRI and rsfMRI-based ADA, including data preparation (e.g., removing backgrounds and data normalisation), 3D MRI region definition, whole-brain matter similarity computation based on pre-defined 3D regions, spectral domain feature extraction, Wilcoxon rank-sum test-based feature selection, and naive Bayers-based depression classification.
4.3 Face/head-based depression assessment
@@
Since human facial/head behaviours provide informative depression biomarkers (discussed in Sec. 2.2.4) while recent advanced DL models can accurately and automatically recognise various facial/head behaviours (e.g., AUs, head movements and facial expressions), automatic visual behaviour-based depression assessment is becoming a prevalent research topic in the past decade. Existing facial/head behaviour-based ADA approaches can be categorised into three types: (i) face image-based approaches which assess depression merely based on a static facial display (e.g., a video frame) [59, 60]; (ii) video-based approaches which assess depression based on spatio-temporal facial behaviours [28, 43, 172, 191, 196, 192]; and (iii) facial attribute-based approaches which assess depression based on pre-extracted anonymous facial descriptors (e.g., facial landmarks [35, 197], AUs [214, 42, 167], head poses [61, 62], gazes/blinks [175, 176] or frame-level deep facial features [63]).
4.3.1 Challenges and solutions
@@
(i) a key challenge for face-based ADA is that depression is better reflected by spatio-temporal facial behaviours rather than static images, i.e., depressed and non-depressed individuals can express same facial displays in a single frame. However, some ADA approaches still model frame-level depression-related facial cues despite the availability of face videos, which frequently aggregate frame-level predictions via average or majority voting for video-level assessment. For example, the baselines and participants of AVEC 2013 [59] and AVEC 2014 challenges [60], generally extract different hand-crafted image-level descriptors (e.g., Histogram of oriented gradients (HOG) [278] and Local Phase Quantisation (LPQ) [279]), and then feed them to SVM Regression (SVR) for assessing frame-level depression severity. Following a similar pipeline, several early studies also extract Gabor wavelet features [39, 250] from static face images, with KNN as the depression predictor. However, such approaches re-use the video-level label as the frame-level labels to train models, and thus may cause ‘ill-posed problems’ (i.e., similar static facial expression inputs paired with different depression labels), making it impossible to learn good hypothesis. To address this issue, spatio-temporal modelling strategies have been particularly explored for ADA (please refer to Sec. 4.3.3 for details); (ii) while depression is associated with long-term and multi-scale facial behaviours [42], most existing spatio-temporal ADA methods only learn depression cues from short segments, ignoring crucial long-term (i.e., video-level) but variable-length spatio-temporal depression cues at the feature extraction level [204, 203]. This is partially addressed by video-level modelling ADA approaches. Song et al. [42, 214] propose a spectral encoding strategy which treats the facial primitives as a multi-channel time-series, applying Fourier Transform to generate a raw spectral representation, and then normalises its size through frequency alignment to standardise all frame-level features as a fixed-size video-level representation. Building on this, Chen et al. [211] extend the method by incorporating a Neural Architecture Search (NAS) in order to identify optimal networks for inferring depression status from the video-level facial primitive spectral representations. Another study [205] investigates a sequential chained-fusion strategy to extract complementary depression patterns contained in both facial display and facial dynamics for the video-level depression assessment. Melo et al. [43] propose a maximization branch to learn multi-scale smooth facial variations and a difference branch for encoding multi-scale facial dynamics from the video-level image sequence, both of which are fused for video-level ADA; (iii) the third limitation is that most existing ADA methods were devoted to assessing adults’ depression status. To bridge this gap, Abbasi et al. [34] construct a video-level spatio-temporal graph for assessing children’s depression status; (iv) another challenge for inferring depression from facial/head behaviours is that these behavuours are more prone to data privacy and ethical issues. Thus, some ADA studies [42, 197, 61, 59] were developed to infer depression from anonymous facial attributes (discussed in Sec. 4.3.4). Alternatively, Pan et al. [212] design a privacy preservation ADA system which records de-identified facial images, while retaining not only depression-relevant features but also emotion-related cues for face image-level ADA; and (v) face/head data can also suffer from noisy and ambiguous depression annotations. He et al. [255] proposes three modules (self-attention, square ranking regularization and re-label modules) to reduce the impact of noisy and ambiguous depression annotations. To deal with missing faces, face detection errors, frame loss and incorrect labels, Song et al. [195] introduces a generic loss function relaxation strategy to mitigate abnormal loss values for various loss functions during video-based ADA model training. We discuss other existing face/head behaviour-based ADA approaches in the following sub-paragraphs.
4.3.2 Face image-based approaches
@@
while depressed and non-depressed individuals can express same static facial displays (e.g., neutral face), some approaches still attempt to assess depression from a single face image (e.g., facial expression and pupil [190, 182]). Melo et al. [49] propose a distribution learning strategy to train a ResNet for face image-based ADA. Liu et al. [198] pre-train a CNN network on a facial expression dataset and then linearly superimpose the entropy of the detected expression into depression severity. Zhao et al. [199] propose a frequency attention strategy to identify depression from static facial displays. Li et al. [182] integrate a dual-scale convolution and a channel-wise adaptive attention to a standard CNN network, aiming to infer depression from a facial expression. Zhang et al. [213] propose a two-stream deep network LKCT. A decoupled knowledge distillation module then distills the LKCT into the lightweight network for face image-based ADA. Similarly, a standard frame-level DL-based ADA pipeline also learns depression cues from individual face frames and aggregates predictions or features for video-level assessment. Zhou et al. [201] propose a DepressNet consisting of two convolution layers and four bottleneck residual convolution blocks to identify depression-informative facial regions from each face frame for ADA. Shang et al. [41] propose to extract each facial image with local textures in the quaternion domain to preserve the dependence of color channels, while deep learning global facial features via a simple CNN. Then, it fuses hand-crafted local features with the deep-learned global cues for frame-level depression assessment. To model multi-scale spatial cues, He et al. [254] propose local and global attention blocks to capture depression-related facial features at multiple spatial scales, which are then combined via a Weighted Spatial Pyramid Pooling for image-level ADA.
4.3.3 Face video-based approaches
@@
Facial temporal dynamics contain crucial biomarkers (discussed in Sec. 2.2.4) for inferring depression. Early face video-based ADA approaches frequently summarise video-level spatio-temporal facial behaviours by computing either the video-level statistics of hand-crafted frame/segment-level facial features (e.g., the major facial expressions, variance of facial expressions [280]) or differences/dynamics between facial frames such as velocity of angular [118], facial appearance eigenviectors [117], three orthogonal planes-based dynamic representations (e.g., LBP-TOP, LGBP-TOP, LPQ-TOP) [115, 251, 24, 187, 193], Motion History Histogram (MHH) [252], and pixel differences between adjacent frames [253]. For instance, Pampouchidou et al. [238] extract three types of features, including appearance-based features (LBP and LBP-TOP), model-based features (active appearance models (AAMs)), and motion-based features (i.e., frame differencing, feature point tracking and optical flow), from every frame/short segment of the given face video, and then feed their statistics to a SVR for ADA. He et al. [187] extend the LBP-TOP feature to MRLBP-TOP for extracting short-term dynamics and then apply Fisher Vector to aggregate them for video-level ADA. SVR then integrates decisions across dimensions for video-level ADA. Instead of directly learning depression cues from face videos, Casado et al.[227] calculate over 60 video-level statistical, geometrical, and physiological features from the remote photoplethysmography signals obtained from each face video for ADA. Yang et al. [173] proposes a Trial Selection Tensor Canonical Correlation Analysis to fuse spatio-temporal relevance between facial dynamic features and pupil diameter features for ADA due to their complementary affective nature.
@@
A prototypical DL solution for face video-based ADA is through learning frame-level facial features/predictions and aggregating them as a video-level depression representation. Uddin et al. [202] apply LSTM to model the facial dynamics from CNN-learned frame-level depression features. Xu et al. [28] first encode a multi-scale facial spatial representation from every frame of the target video, and then integrate all frame-level multi-scale spatial features as a single spectral graph representation for video-level ADA in an effort to learn a richer and more condensed latent representation. Pan et al. [203] integrate both spatial and temporal attention mechanisms to 3D CNNs, and then combine all segment-level depression predictions as the video-level prediction. Similarly, Al et al. [204] utilise a C3D network to extract short segment-level facial dynamics and then averages all segment-level predictions for video-level ADA. Li et al. [183] and Melo et al. [209] propose two-stream networks that jointly learns facial image-level depression cues and video-level facial dynamics, both of which are concatenated via a Multi-Layer Perceptron (MLP) for the final video-level ADA. Zhu et al. [206] propose two-stream networks that jointly learns facial appearance features and short-term facial dynamics via optical flow. Zhang et al. [207] propose a lightweight CNN incorporating temporal difference for short-term facial behaviours, followed by an attention for long-term facial behaviour modelling, which are fused via Interactive Multi-head Attention for video-level ADA. He et al. [208] employ a video Swin-Transformer to learn spatio-temporal and fine-grained depression-specific features, integrating them via a Depression Channel Attention Fusion (DCAF) module for video-level ADA. Xie et al. [184] design an end-to-end ADA framework for variable-length videos, which leverages a 3D CNN for local temporal pattern extraction and a Redundancy-Aware Self-Attention (RAS) scheme for global feature aggregation. Moreno et al. [210] pre-train a CNN on action recognition datasets and fine-tune it on ADA face videos, enabling analysis of frame-level regions and short-term temporal expressions.
4.3.4 Facial behaviour primitive-based approaches